The half life of neutron is 693 seconds.What fraction of neutrons will decay when a beam of neutrons having kinetic energy 0.084 eV travells a distance of 1km? (mass of neutron =1.68times10^(-27)kg and ln2=0.693)
Ans 25×10^-5
1
Expert's answer
2020-01-23T06:09:15-0500

"T=693 \; s"

"KE=0.084 \;eV=0.084 \times 1.6\times 10^{-19} \; J"

"s=1000 \;m"

"m=1.68\times 10^{-27} \; kg"

"\ln 2=0.693"


Solution:

"KE=\frac{mv^2}{2}"

"v=\sqrt {\frac{2KE}{m}} =\sqrt {\frac{2\times0.084\times1.6\times 10^{-19}}{1.68\times 10^{-27}}}=4\times 10^3 \; m/s"


Time to travel distance "s=1000 \;m:"

"t=\frac{s}{v}=\frac{1000}{4\times 10^3}=0.25 \;s"


Exponential decay equation:

"N(t)=N_02^{\frac{-t}{T}}"

"t=0.25\; s: \;"

"\frac{N}{N_0}=2^\frac{-0.25}{693}=2^{-3.607\times 10^{-4}}=0.99975" - fraction of neutrons that won’t decay

"1-\frac{N}{N_0} =1-0.99975=0.00025" - fraction of neutrons that will decay


Answer: "25\times 10^{-5}"